OpenStudy (anonymous):

Relationship B has a lesser rate than Relationship A. The graph represents Relationship A. Which equations could represent Relationship B? Choose ALL answers that are correct. A. y=4x B. y=x C. y=3/2x

OpenStudy (anonymous):

@texaschic101

OpenStudy (igreen):

First find the slope of Relationship A. \(m = \dfrac{y_2-y_1}{x_2-x_1}\) Plug in the two points: \(m = \dfrac{6-3}{2-1}\) Subtract: \(m = \dfrac{3}{1}\) Simplify: \(m = 3\) So the slope of Relationship A is 3. Relationship B's slope is LESS than Relationship A's slope. So Relationship B's slope is less than 3..so which answers are correct? @Samus_Aran

OpenStudy (igreen):

Where's Option D? @Samus_Aran

OpenStudy (anonymous):

oh sorry I was afk having lunch. D. y=5x

OpenStudy (anonymous):

@iGreen

OpenStudy (igreen):

Okay, which options have a slope less than 4?

OpenStudy (anonymous):

B. y=x C.=3/2x

OpenStudy (igreen):

Yep, so those are your answers.

OpenStudy (anonymous):

thank you :D

OpenStudy (anonymous):

100% :D